OpenStudy (anonymous):

@ganeshie8 @campbell_st @BSwan @kohai

OpenStudy (anonymous):

Write as an exponent: log2 1/2=-1

OpenStudy (ikram002p):

you mean \(log_2 (1/2 )=-1 \) take 2 power for both sides 1/2 =2^-1

OpenStudy (anonymous):

is that the answer